OpenStudy (anonymous):

solve the equation. 5(3-4x)=7-(4-x)

OpenStudy (anonymous):

x=4/7

OpenStudy (anonymous):

@mitchelsewbaran how did you get that?

OpenStudy (anonymous):

5(3-4x)=7-(4-x) 15-20x=7-(4-x) 15-20x=7+x-4 15-20x=3+x 15-20x-x=3 15-21x=3 -21x=-15+3 -21x=-12 x= 12/-21 Answ: x= 4/7